Would someone please help me with this one:
Solve the following quadratic equations by factoring or by any of the quadratic equations available such as square root, completing the square, etc.
4x^2-8x+3=5

Answer by Nate(3500)

Firstly, it is best to try and factor it:
4x^2 - 8x + 3 = 5
4x^2 - 8x - 2 = 0
It can't be factored, so the next simple thing would be to complete the square.
4x^2 - 8x = 2
x^2 - 2x = 1/2
(x - 1)^2 = 3/2
x - 1 = +- sqrt(1.5)
x = 1 +- sqrt(1.5)